Compare the partial differential equation of Wave and Heat; state their similari ties and differences.
1
Expert's answer
2020-11-17T13:01:39-0500

The wave equation has a form

"\displaystyle \frac{\partial^2u}{\partial t^2} = c^2 \Delta u" (1)

The heat equation is following

"\displaystyle \frac{\partial u}{\partial t} = \alpha \Delta u"

Similarities:

1) both are second-order PDE

2) both have "\Delta u" on the right-hand side with some coefficient

Differences:

1) (1) is hyperbolic equation, (2) is parabolic equation

2) Coefficients have different physical sense and different power: c in (1) is speed of the wave, "\alpha" in (2)is thermal diffusivity

3) they have different order of time derivative, and, therefore, the solutions depend on time in a different way (for (1) it is "\sim e^{-t}", for (2) it is "\sim ct")